OpenStudy (anonymous):

The average annual income, l, in dollars, of a lawyer with an age of x years is modeled with the following function: I = -425x^2 + 45,000x - 650,000 According to this model, what is the maximum average annual income, in dollars, a lawyer can earn? Round to the nearest whole dollar. okay, i can't figure this out. every time i plug it in to the quadratic formula i get a low number like 90 or 16. i'm terrible with word problems, can someone try to walk me through this?

OpenStudy (anonymous):

i tried it twice more and this time got 69 and then 354 so obviously i don't know what i'm doing :(

OpenStudy (anonymous):

Hey, I'm sorry but I'm going right now (for a while), maybe if you're online later, I can help you. But maybe someone will help you by the time I come back online, take care and good luck. I will get back to you later.

OpenStudy (anonymous):

@iambatman okay, sounds good. thanks anyways!

OpenStudy (solomonzelman):

The question is asking for annual income, and thus that means that x=1. In your question, the equation for income is \(\normalsize\color{red}{ \rm I = -425x^2 + 45,000x - 650,000 }\) And x is the number of years. So for annual income, x=1. Plug that in and find y. \(\normalsize\color{blue}{ \rm I = -425x^2 + 45,000x - 650,000 }\) \(\normalsize\color{blue}{ \rm I = -425~\color{orangered}{(1)}^2 + 45,000~\color{orangered}{(1)} - 650,000 }\) \(\normalsize\color{blue}{ \rm I = -425 + 45,000 - 650,000 }\) \(\normalsize\color{blue}{ \rm I = -605425 }\)

OpenStudy (solomonzelman):

Something that really seems to be the correct solution comes out to make no sense what so ever -:(

OpenStudy (anonymous):

@SolomonZelman thank you so much anyway, according to the work my teacher sent me i was supposed to graph the function and find it's vertex, but that wasn't how i was taught so i'm still a little confused

jimthompson5910 (jim_thompson5910):

-425x^2 + 45,000x - 650,000 is in the form ax^2 + bx + c where a = -425 b = 45000 c = -650000

jimthompson5910 (jim_thompson5910):

the x coordinate of the vertex is found using this formula x = -b/(2a) so plug in a = -425 and b = 45000 to get x = -b/(2a) x = -(45000)/(2(-425)) x = 52.941176 <--- this is approximate Once you have the x coordinate of the vertex, you plug this x value back into y = -425x^2 + 45,000x - 650,000 to get the y coordinate of the vertex. I'll let you do this part.

OpenStudy (anonymous):

@jim_thompson5910 I got the y-coordinate as 541,176.47 is that wrong?

jimthompson5910 (jim_thompson5910):

I'm getting the same as well